OpenStudy (anonymous):

Find the area under the graph of f(x) = e-3ln(x) on the interval [1, 2].

jimthompson5910 (jim_thompson5910):

is the function \[\LARGE f(x) = e^{-3\ln(x)}\] ??

OpenStudy (anonymous):

Yes

OpenStudy (jhannybean):

Then what we're going to do is integrate the function from 1 to 2

OpenStudy (freckles):

you can make that function prettier before doing that whole integration thingy

jimthompson5910 (jim_thompson5910):

You can simplify things a bit first \[\LARGE f(x) = e^{-3\ln(x)}\] \[\LARGE f(x) = e^{\ln(x^{-3})}\] \[\LARGE f(x) = x^{-3}\]

OpenStudy (jhannybean):

\[\huge e^{-3\ln_e (x)} = e^{\ln_e(x^{-3})} = x^{-3} \] Does that make sense to you though?

OpenStudy (anonymous):

I got 0.375

OpenStudy (anonymous):

Yeah I understand how it's done.

OpenStudy (jhannybean):

\[\large \int_1^2 x^{-3}dx = \left.\frac{x^{-3+1}}{-3+1}\right]_{1}^2 = \left.\frac{x^{-2}}{-2}\right]_1^2=-\frac{1}{2}\left( \frac{1}{4} - 1\right) = \frac{3}{8} \qquad \checkmark\]